# On multiplicities in length spectra of semi-arithmetic hyperbolic surfaces

### v1

@@ 1. Introduction
Let $\𝑆\ΓS=\Gamma\backslash\mathbb{H}$ be a hyperbolic surface, possibly with elliptic singularities and with cusps. The prime geodesic theorem, which goes back to the classical work of Hubert and Selberg, asserts that the number of closed geodesics on $S$ of length at most $\ell$ grows exponentially:
$\mathcal{N}(\ell)\sim\dfrac{e^{\ell}}{\ell}\text{ as }\ell\to\infty.$
Distribution of the lengths of closed geodesics (periodic orbits) on surfaces has been of interest for theoretical physics in connection with the phenomenon of quantum chaos. In 1980s, Aurich and Steiner experimentally observed that some surfaces have extremely large multiplicities in their geodesic length spectrum, more precisely, they detected the strong exponential growth of the mean multiplicities (EGMM) in the spectrum of these particular surfaces [AS88]:
$\langle g(\ell)\rangle=\dfrac{\mathcal{N}(\ell)}{\mathcal{N}^{\prime}(\ell)}\sim c\dfrac{e^{\ell/2}}{\ell}\text{ as }\ell\to\infty,$
where $\mathcal{N}^{\prime}(\ell)$ denotes the number of distinct lengths of geodesics in the interval $(0,\ell]$. It was later noticed in [BGGS92] that EGMM tends to occur in the spectrum of arithmetic surfaces, i.e. when $\Gamma$ is an arithmetic Fuchsian group. A short proof of this fact was given by Luo and Sarnak in [LS94] (see also [Bol93a] for a detailed derivation which in addition yields the value for the multiplicative constant in the asymptotic growth formula).
@@
where $\mathcal{L}(\Gamma)=\{|\mathrm{tr}(\gamma)|\mid\gamma\in\Gamma\}$. The absolute values of traces of hyperbolic isometries determine the lengths of corresponding geodesics on the quotient surface (see Section 2), hence we defined the bounded clustering in the length spectrum of $S$. Sarnak conjectured that the B-C property is characteristic for arithmetic Fuchsian groups, which was later proved in the non-cocompact case by Geninska and Leuzinger [GL08]. Around the time of Sarnak’s conjecture it was assumed and stated in some papers that the EGMM property should also imply arithmeticity (see e.g. [Bol93b]). Later on, however, physicists found experimental evidence and heuristic arguments showing that some non-arithmetic hyperbolic surfaces also tend to have EGMM, although with a smaller exponent [BGGS97, BS04]. In this paper, we prove this fact. To our best knowledge, this is the first time when exponential growth of the mean multiplicities in the length spectrum is rigorously confirmed in the non-arithmetic case.
@@
Our main results are given in Theorems 3.4 and 3.7 below that state that all cocompact (resp., non-cocompact) semi-arithmetic Fuchsian groups which admit a modular embedding and have arithmetic dimension at most $2$ have exponential growth of the mean multiplicity in the length spectrum. We will give definitions of (semi-)arithmeticity, modular embedding and arithmetic dimension in Section 2. The proof of the first theorem uses the Schwarz–Pick lemma and some arithmetic properties of semi-arithmetic groups. The latter imply a weaker form of the B-C property which is sufficient for our purpose (see Lemma 3.2). This argument does not generalize directly to the non-compact surfaces. In order to prove Theorem 3.7 we first apply the generalized Kirszbraun theorem of [LS97], which allows us to work with the compact core of the surface $S$. With this we are able to replicate the proof of Theorem 3.4. The details are given in Section 3.
@@
For arithmetic dimension $1$ our theorems refer to the well known case of the arithmetic Fuchsian groups, the new result is about groups whose arithmetic dimension is equal to $2$. In Section 4 we list examples of such groups. By Cohen and Wolfart [CW90], all Fuchsian triangle groups admit modular embeddings. Nugent and Voight showed in [NV17] that there exist only finitely many triangle groups of a given arithmetic dimension (see also [BCDTP25] for a different proof of a more general result). The method of [NV17] is effective and allows to produce a list of such groups, for small dimensions the results are presented in Section 6 of their paper. The list of triangle groups of arithmetic dimension $2$ consists of $148$ cocompact and $16$ non-cocompact groups. The cocompact part of the list includes four groups from the Hilbert series (cf. [McM23]). The non-cocompact part includes a remarkable non-arithmetic Hecke group $\Delta(2,5,\infty)$ which was closely studied in a recent paper by McMullen [McM22]. The previous work of McMullen also provides an infinite set of examples of non-cocompact semi-arithmetic Fuchsian groups of arithmetic dimension $2$ which admit a modular embedding and are not commensurable with the triangle groups [McM03].
@@
Acknowledgments. We thank Jens Marklof and Curtis McMullen for their interest in this work and comments. The work of M.B. is partially supported by the FAPERJ grant E-26/204.250/2024. G.C. is grateful for the grant 2024/01650-3, São Paulo Research Foundation (FAPESP). The authors C.D. and G.T.P. are grateful for the support of CNPq Grant 408834/2023-4.
@@ 2. Preliminaries
This is a natural generalization of arithmetic Fuchsian groups. Indeed, by using a characterization given by Takeuchi, arithmetic Fuchsian groups must also satisfy the extra condition that every non-trivial Galois embedding $K\hookrightarrow\mathbb{R}$ maps the set of traces of elements of $\Gamma^{(2)}$ to a bounded subset $[-2,2]\subset\mathbb{R}$ (cf. [Tak75]). In the semi-arithmetic case we allow $r\geq 1$ embeddings of $K$ for which the images of the set of traces are unbounded. The number $r$ is called the arithmetic dimension of $\Gamma$ (cf. [NV17]). Notice that arithmetic Fuchsian groups have arithmetic dimension $1$, and in general we have $1\leq r\leq d$, with $d=[K:\mathbb{Q}]$ denoting the degree of the invariant trace field.
@@
Let $\sigma_{1}=\mathrm{id},\sigma_{2},\ldots,\sigma_{d}$ be the Galois embeddings of $K$ into $\mathbb{R}$. We assume that the associated quaternion algebra $A/K$ splits over $\sigma_{1}=\mathrm{id}$, …, $\sigma_{r}$. Thus we have an isomorphism
@@
From now on we restrict to the case $r=2$, which is our primary interest in this paper.
@@
Let $\Gamma$ be a semi-arithmetic Fuchsian group, with arithmetic dimension 2. The group $\Gamma^{(2)}$ is derived from a quaternion algebra and has a family of embeddings $f:\Gamma^{(2)}\to\mathrm{SL}(2,\mathbb{R})^{2}$ given by restriction of $\rho\circ\rho_{1}^{-1}$ for each choice of $\rho=(\rho_{1},\rho_{2})$ as above. We say that $\Gamma$ admits a modular embedding if $f$ can be chosen so that there exists an $f$-equivariant holomorphic (or anti-holomorphic) function $\tilde{F}:\mathbb{H}\to\pm\mathbb{H}\times\pm\mathbb{H}$, that is, a function $\tilde{F}=(\tilde{F}_{1},\tilde{F}_{2})$ satisfying
@@
We conclude this section with some geometric properties of semi-arithmetic groups. Let $\Gamma$ be a semi-arithmetic group of arithmetic dimension $2$, $\Delta<\mathrm{SL}(2,\mathbb{R})^{2}$ its ambient arithmetic group, and
$S=\Gamma\backslash\mathbb{H}\hookrightarrow X=\Delta\backslash(\pm\mathbb{H}\times\pm\mathbb{H})$
@@
    If we consider the natural Riemannian metric on $-\mathbb{H}$ which makes complex conjugation an isometry, the surface $S$ is not totally geodesic in $X$ with the locally symmetric metric induced by the projection (otherwise, the subgroup $\Gamma$ would be arithmetic by Bergeron–Clozel [BC05, Proposition 15.2.2] and the remark immediately after the proof of the proposition addressing the non-cocompact case).
@@
    The group $\Gamma$ is Zariski dense in $G=\mathrm{SL}(2,\mathbb{R})^{2}$ (see [McM23, Section 5]).
@@ 3. Results
Let $\Gamma$ be a cocompact semi-arithmetic Fuchsian group of arithmetic dimension $2$ which admits a modular embedding, e.g. $\Gamma=\Delta(2,6,10)$ (see [McM23]). In what follows, we denote by $K$ the invariant trace field of $\Gamma$ and by $\mathcal{O}$ the ring of integers of $K$. The field $K$ is a totally real number field of degree $d\geq 2$ whose Galois embeddings into $\mathbb{R}$ will be denoted by $\sigma_{1}=\mathrm{id},\sigma_{2}=\sigma,\sigma_{3},\ldots,\sigma_{d}$.
@@
for all $\gamma\in\Gamma^{(2)}$, where $\gamma^{\sigma}$ denotes $\rho_{2}\circ\rho_{1}^{-1}(\gamma)$. The notation is motivated by the fact that the trace and norm of $\gamma^{\sigma}$ are equal to the trace and norm of the $\sigma$-conjugate of $\gamma$.
@@
Let $\Gamma$ be a cocompact semi-arithmetic Fuchsian group of arithmetic dimension $2$ which admits a modular embedding. Let $\sigma:K\to\mathbb{R}$ be the unique non-trivial Galois embedding with $\sigma(\mathrm{tr}(\Gamma^{(2)}))\not\subset[-2,2]$. Then there exists $10<\delta=\delta(\Gamma)<$ 1 such that
@@
We can assume that $\Gamma$ is derived from a quaternion algebra. Let $\Omega$ be a fundamental domain for $\Gamma$ contained in a compact set. Since the map $F$ defined above is holomorphic (or anti-holomorphic), by the Schwarz–Pick lemma, there exists $10<\delta<$ 1 such that
@@
If $\gamma^{\sigma}$ is hyperbolic with displacement $\ell^{\prime}$, we have
@@
[…]
@@
Let $\Gamma$ be as above. We define the counting function of traces:
$\mathcal{L}^{\prime}(T)=\#\{t\in(2,T]:t=|\mathrm{tr}(\gamma)|\text{ for some }\gamma\in\Gamma\}.$
@@
Let $\Gamma$ be a cocompact semi-arithmetic group satisfying assumptions of Lemma 3.1. Then there exist $10<\delta<$ 1 and $C>0$ depending on $\Gamma$ such that for any $N>2$ we have
#$\#(\mathcal{L}(\Gamma)\cap[N-1,N])\leq CN^{1-\delta}.$
Moreover, if $\Gamma$ is derived from a quaternion algebra, then any two different traces in $\mathcal{L}^{\prime}(T)$ are $cT^{\delta-1}$ separated for a constant $c>0$ depending only on the degree of $K$.
@@
First assume that $\Gamma$ is derived from a quaternion algebra. Consider the set $\{\mathrm{id},\sigma,\sigma_{3},\ldots,\sigma_{d}\}$ of the Galois embeddings of $K$. We recall that $|\sigma_{j}(\mathrm{tr}(\gamma))|<2$ for any hyperbolic element $\gamma\in\Gamma^{(2)}$ and for all $j=3,\ldots,d$. By inequality (4) from Lemma 3.1, if $t,s\in(2,T]$ are different traces of elements of $\Gamma$, then
@@
for some constant $C_{1}>0$ which depends only on the degree of $K$. This inequality shows that $t$ and $s$ are $cT^{\delta-1}$ separated with $c=\frac{1}{C_{1}}>0.$
In general, similar to [LS94, Lemma 2.1], we consider all the embeddings of the field $L=\mathbb{Q}(\mathcal{L}(\Gamma))$ into $\mathbb{C}$ which we denote by $\tau_{1}=\mathrm{id},\tau_{2},\ldots,\tau_{n}$, $n=[L:\mathbb{Q}]$. We order them so that the first $k$ embeddings extend the embedding $\sigma_{1}=\mathrm{id}$ of the subfield $K$ of $L$, and the next $k$ embeddings extend $\sigma_{2}=\sigma$. Since $\mathrm{tr}(\gamma)^{2}=\mathrm{tr}(\gamma^{2})+2$ we have $x^{2}-2\in\mathcal{L}(\Gamma^{(2)})$ for any $x\in\mathcal{L}(\Gamma)$, in particular, $x^{2}\in K$. Suppose there are more than $2^{2k-1}(M+1)$ points of $\mathcal{L}(\Gamma)$ in the interval $[N-1,N]$, with $M$ a positive integer. Then by the pigeonhole principle there are at least two different elements $t,s\in\mathcal{L}(\Gamma)$ and some $\epsilon_{i}\in\{-1,1\}$, $i=2,\ldots,2k$, such that $0<|t-s|\leq\frac{1}{M}$, $\tau_{i}(t)=\epsilon_{i}t$, $\tau_{i}(s)=\epsilon_{i}s$ for $2\leq i\leq k,$ and $\tau_{i}(t)=\epsilon_{i}\sqrt{\sigma_{2}(t^{2})}$, $\tau_{i}(s)=\epsilon_{i}\sqrt{\sigma_{2}(s^{2})}$ for $k<i\leq 2k$. Thus $0\neq t-s\in\mathcal{O}_{L}$, $|\tau_{i}(t-s)|\leq\frac{1}{M}$, $1\leq i\leq k$, by the assumption, $|\tau_{i}(t-s)|\leq 4N^{1-\delta}$, $k<i\leq 2k$, by Lemma 3.1, and $|\tau_{i}(t-s)|\leq 4$ for $i>2k$ (we assume that $\tau_{i}(\mathcal{L}(\Gamma))\subset[-2,2]$ for $i>2k$, which follows from $\mathrm{tr}(\gamma)^{2}=\mathrm{tr}(\gamma^{2})+2$ and $\sigma_{j}(\mathcal{L}(\Gamma^{(2)}))\subset[-2,2]$, for $j>2$). Hence we have
$1\leq|\mathrm{N}_{L\mid\mathbb{Q}}(t-s)|\leq C_{2}\frac{1}{M^{k}}(4N)^{(1-\delta)k}4^{n-2k},$
which implies $M\leq CN^{1-\delta}$ with $C=C(\Gamma)>0$.
@@
There exist $10<\delta<$ 1 and $C>0$ depending on $\Gamma$ such that for $T>2$,
$\mathcal{L}^{\prime}(T)\leq CT^{2-\delta}.$
@@
The result if already known for arithmetic groups by [LS94], so we can assume that $\Gamma$ is a non-arithmetic group as above.
Let $\mathcal{N}^{\prime}(\ell)$ be the number of distinct lengths bounded by $\ell$ of closed geodesics in the orbifold $\.\Γ\Gamma\backslash\mathbb{H}.$ Each length produces a unique positive trace of a hyperbolic element of $\Gamma$ bounded by $2\operatorname{arcosh}({\ell}/{2})$, i.e. with the notation of the previous lemma, we have
(7) $\mathcal{N}^{\prime}(\ell)\leq\mathcal{L}^{\prime}\left(2\cosh\left(\frac{\ell}{2}\right)\right).$
Thus, by Corollary 3.3 and the well known inequality $\cosh(\frac{\ell}{2})\leq e^{\frac{\ell}{2}}$, already used in the proof of Lemma 3.1, we obtain
@@
for some constants $C^{\prime}>0$ and $0<\delta^{\prime}<\frac{1}{2}.$ By the prime geodesic theorem, the number $\mathcal{N}(\ell)$ of closed geodesics of $\Γ\Gamma\backslash\mathbb{H}$ counted with multiplicity is given by
[…]
Thus the mean multiplicity $\langle g(\ell)\rangle$ of closed geodesics of length at most $\ell$ grows at least as
$\langle g(\ell)\rangle=\dfrac{\mathcal{N}(\ell)}{\mathcal{N}^{\prime}(\ell)}\gtrsim c\dfrac{e^{\delta^{\prime}\ell}}{\ell},$
where $c>0$ and $0<\delta^{\prime}<\frac{1}{2}$ depend on $\Gamma$ and its arithmetic invariants. ∎
@@
Since $X$ has curvature $\leq\kappa$ for some $\kappa\leq 0$ and $S$ has curvature $\leq-1$ in the sense of Alexandrov, we can apply the generalized Kirszbraun theorem [LS97, Theorem A] in order to obtain an extension $g$ of $f\mid\Omega$ with the same Lipschitz constant.
@@
Let $\Gamma$ be a non-cocompcat semi-arithmetic Fuchsian group of arithmetic dimension $2$ which admits a modular embedding. Then there exist constants $10<\delta<$ 1 and $C>0$ depending on $\Gamma$ such that
$|\sigma(\mathrm{tr}(\gamma))|\leq C|\mathrm{tr}(\gamma)|^{1-\delta}$
@@
Let us first assume that the group $\Gamma$ is torsion-free and derived from a quaternion algebra. Consider the hyperbolic surface $\𝑆\ΓS=\Gamma\backslash\mathbb{H}$ and a lattice $\tilde{\Gamma}<\mathrm{SL}(2,\mathbb{R})\times\mathrm{SL}(2,\mathbb{R})$ with $\rho(\gamma):=(\gamma,\gamma^{\sigma})\in\tilde{\Gamma}$ for all $\gamma\in\Gamma.$
Let $\Omega\subset S$ be a compact core of $S$. If $F:\mathbb{H}\to\pm\mathbb{H}$ is a holomorphic or anti-holomorphic map such that $F(\gamma\cdot z)=\gamma^{\sigma}\cdot F(z)$, for any $z\in\mathbb{H}$ and ${}\gamma\in\Gamma$, provided by the modular embedding of $\Gamma$. Consider the metric space $X=\tilde{\Gamma}\backslash(\pm\mathbb{H}\times\pm\mathbb{H})$, which is locally isometric to $\mathbb{H}\times\mathbb{H}$ equipped with the product metric. We have an embedding $G:S\to X,$ given by
$G(\Gamma z)=\tilde{\Gamma}(z,F(z)).$

### v2

@@ 1. Introduction
[…]
Distribution and multiplicity of the lengths of closed geodesics (periodic orbits) on surfaces have been of interest for theoretical physics in connection with the phenomenon of quantum chaos. In 1980s, Aurich and Steiner experimentally observed that some surfaces have extremely large multiplicities in their geodesic length spectrum. This result is essentially the starting point for our present work so let us give some definitions and state it precisely.
Let $\{0<\ell_{1}<\ell_{2}<\cdots\}$ be the geodesic length set (that is, without multiplicities) of $S$ with counting function
$\mathcal{N}^{\prime}(\ell)=\#\{n\,;\,\ell_{n}\leq\ell\}.$
Denote by $g(\ell_{n})$ the multiplicity of $\ell_{n}$ in the geodesic length spectrum of $S$. The mean multiplicity $\langle g(\ell)\rangle$ is defined to be a continuous function of $\ell$ that is an average order of $g$, meaning:
$\mathcal{N}(\ell)=\sum_{\ell_{n}\leq\ell}g(\ell_{n})\sim\sum_{\ell_{n}\leq\ell}\langle g(\ell_{n})\rangle=\int_{0}^{\ell}\langle g(x)\rangle\,d\mathcal{N^{\prime}}(x),$
as $\ell\to\infty$. The right-hand side is the Riemann–Stieltjes integral of the mean multiplicity with respect to $\mathcal{N^{\prime}}$.
We refer to [Bol93b, Section 3.2] for more details regarding this notion. We will say that $S$ has exponential growth of the mean multiplicities (EGMM) if for sufficiently large $\ell$ its mean multiplicity satisfies
$\langle g(\ell)\rangle\geq e^{C\ell}\text{ for some constant }C=C(S)>0.$
[…]
@@
where $\mathcal{L}(\Gamma)=\{|\mathrm{tr}(\gamma)|\mid\gamma\in\Gamma\}$. The absolute values of traces of hyperbolic isometries determine the lengths of corresponding geodesics on the quotient surface (see Section 2), hence we defined the bounded clustering in the length spectrum of $S$. Sarnak conjectured that the B-C property is characteristic for arithmetic Fuchsian groups, which was later proved in the non-cocompact case by Geninska and Leuzinger [GL08]. Around the time of Sarnak’s conjecture it was assumed and stated in some papers that the EGMM property should also imply arithmeticity (see e.g. [Bol93b]). Later on, however, physicists found experimental evidence and heuristic arguments showing that some non-arithmetic hyperbolic surfaces also tend to have EGMM, although with a smaller exponent [BGGS97, BS04]. In this paper, we prove this fact. To the best of our knowledge, this is the first time when exponential growth of the mean multiplicities in the length spectrum is rigorously confirmed in the non-arithmetic case.
Our main result is the following
Theorem 1.1.
All semi-arithmetic Fuchsian groups which admit a modular embedding and have arithmetic dimension at most $2$ have exponential growth of the mean multiplicity in the length spectrum.
@@
We will give definitions of (semi-)arithmeticity, modular embedding and arithmetic dimension in Section 2. The proof is different for cocompact and non-cocompact Fuchsian groups, which are treated in Theorem 3.4 and Theorem 3.7, respectively. The proof of the first theorem uses the Schwarz–Pick lemma and some arithmetic properties of semi-arithmetic groups. The latter imply a weaker form of the B-C property which is sufficient for our purpose (see Lemma 3.2). This argument does not generalize directly to the non-compact surfaces. In order to prove Theorem 3.7 we first apply the generalized Kirszbraun theorem of [LS97], which allows us to work with the compact core of the surface $S$. With this we are able to replicate the proof of Theorem 3.4. The details are given in Section 3.
@@
For arithmetic dimension $1$ our theorem refers to the well known case of the arithmetic Fuchsian groups, the new result is about groups whose arithmetic dimension is equal to $2$. In Section 4 we list examples of such groups. By Cohen and Wolfart [CW90], all Fuchsian triangle groups admit modular embeddings. Nugent and Voight showed in [NV17] that there exist only finitely many triangle groups of a given arithmetic dimension (see also [BCDTP25] for a different proof of a more general result). The method of [NV17] is effective and allows us to produce a list of such groups, for small dimensions the results are presented in Section 6 of their paper. The list of triangle groups of arithmetic dimension $2$ consists of $148$ cocompact and $16$ non-cocompact groups. The cocompact part of the list includes four groups from the Hilbert series (cf. [McM23]). The non-cocompact part includes a remarkable non-arithmetic Hecke group $\Delta(2,5,\infty)$ which was closely studied in a recent paper by McMullen [McM22]. The previous work of McMullen also provides an infinite set of examples of non-cocompact semi-arithmetic Fuchsian groups of arithmetic dimension $2$ which admit a modular embedding and are not commensurable with the triangle groups [McM03].
@@
Acknowledgments. We thank Jens Marklof and Curtis McMullen for their interest in this work. We also thank the anonymous referee for careful proofreading and helpful comments. The work of M.B. is partially supported by the FAPERJ grant E-26/204.250/2024, the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université), and LabEx CARMIN (ANR-10-LABX-59-01). G.C. is grateful for the grant 2024/01650-3, São Paulo Research Foundation (FAPESP). The authors C.D. and G.T.P. are grateful for the support of CNPq Grant 408834/2023-4.
@@ 2. Preliminaries
This is a natural generalization of arithmetic Fuchsian groups. Indeed, by using a characterization given by Takeuchi, arithmetic Fuchsian groups must also satisfy the extra condition that every non-trivial embedding $K\hookrightarrow\mathbb{R}$ maps the set of traces of elements of $\Gamma^{(2)}$ to a bounded subset $2,2]\subset\mathbb{R}$ (cf. [Tak75]). In the semi-arithmetic case we allow $r\geq 1$ embeddings of $K$ for which the images of the set of traces are unbounded. The number $r$ is called the arithmetic dimension of $\Gamma$ (cf. [NV17]). Notice that arithmetic Fuchsian groups have arithmetic dimension $1$, and in general we have $1\leq r\leq d$, with $d=[K:\mathbb{Q}]$ denoting the degree of the invariant trace field.
@@
Let $\sigma_{1}=\mathrm{id},\sigma_{2},\ldots,\sigma_{d}$ be the embeddings of $K$ into $\mathbb{R}$. We assume that the associated quaternion algebra $A/K$ splits over $\sigma_{1}=\mathrm{id}$, …, $\sigma_{r}$. Thus we have an isomorphism
@@
$2.$
From now on we restrict to the case $r=2$, which is our primary interest in this paper. Let $\Gamma$ be a semi-arithmetic Fuchsian group, with arithmetic dimension 2.
Notice that the group $\Gamma^{(2)}$ is derived from a quaternion algebra and has a family of embeddings $f:\Gamma^{(2)}\to\mathrm{SL}(2,\mathbb{R})^{2}$ given by restriction of $\rho\circ\rho_{1}^{-1}$ for each choice of $\rho=(\rho_{1},\rho_{2})$ as above.
@@
We say that $\Gamma$ admits a modular embedding if $f$ can be chosen so that there exists an $f$-equivariant holomorphic (or anti-holomorphic) function $\tilde{F}$$\mathbb{H}\to\pm\mathbb{H}\times\pm\mathbb{H}$, that is, a function $\tilde{F}=(\tilde{F}_{1},\tilde{F}_{2})$ satisfying
@@
We conclude this section with some geometric properties of semi-arithmetic groups. Let $\Delta<\mathrm{SL}(2,\mathbb{R})^{2}$ be the ambient arithmetic group of $\Gamma$ and
[…]
@@
    The group $\Gamma^{(2)}$ is Zariski dense in $\mathrm{SL}(2,\mathbb{R})^{2}$ (see [McM23, Section 5]).
@@
    If we consider the natural Riemannian metric on -$\mathbb{H}$ which makes complex conjugation an isometry, the surface $S$ is not totally geodesic in $X$ with the locally symmetric metric induced by the projection (otherwise, the subgroup $\Gamma$ would be arithmetic by Bergeron–Clozel [BC05, Proposition 15.2.2] and the remark immediately after the proof of the proposition addressing the non-cocompact case). In particular, this shows that the map $F$ is not an automorphism.
@@ 3. Results
Let $\Gamma$ be a cocompact semi-arithmetic Fuchsian group of arithmetic dimension $2$ which admits a modular embedding. In what follows, we denote by $K$ the invariant trace field of $\Gamma$ and by $\mathcal{O}$ the ring of integers of $K$. The field $K$ is a totally real number field of degree $d\geq 2$ whose embeddings into $\mathbb{R}$ will be denoted by $\sigma_{1}=\mathrm{id},\sigma_{2}=\sigma,\sigma_{3},\ldots,\sigma_{d}$.
@@
for all $\gamma\in\Gamma^{(2)}$, where $\gamma^{\sigma}$ denotes $\rho_{2}\circ\rho_{1}^{-1}(\gamma)$. The notation is motivated by the fact that the trace and norm of $\gamma^{\sigma}$ are equal to the trace and norm of the $\sigma$-conjugate of $\gamma$. In particular, $\gamma^{\sigma}$ is parabolic if and only if $\gamma$ is parabolic.
@@
Let $\sigma:K\to\mathbb{R}$ be the unique non-trivial embedding with $\sigma(\mathrm{tr}(\Gamma^{(2)}))\not\subset[-2,2]$. Then there exists $0<\delta=\delta(\Gamma)<1$ such that
@@
We can assume that $\Gamma$ is derived from a quaternion algebra. Let $\Omega$ be a fundamental domain for $\Gamma$ contained in a compact set. The map $F$ defined above is holomorphic (or anti-holomorphic) and not an automorphism (see property (iii) from the previous section or [SSW00, Theorem 3 and Corollary 5]). By the Schwarz–Pick lemma, there exists $0<\delta<1$ such that
@@
Since $\gamma$ is hyperbolic, $\gamma^{\sigma}$ is either hyperbolic or elliptic. If $\gamma^{\sigma}$ is hyperbolic with displacement $\ell^{\prime}$, we have
@@
$=$~ $d(F(z),F(\gamma z))$
$\leq$~ $(1-\delta)\ell$
$=$~ $(1-\delta)2\operatorname{arcosh}\left(\frac{|\mathrm{tr}(\gamma)|}{2}\right).$
@@
There exist $0<\delta<1$ and $C>0$ depending on $\Gamma$ such that for any $N>2$ we have
$\#(\mathcal{L}(\Gamma)\cap[N-1,N])\leq CN^{1-\delta}.$
Moreover, if $\Gamma$ is derived from a quaternion algebra, then any two different traces in $\mathcal{L}(\Gamma)\cap[2,T]$ are $cT^{\delta-1}$ separated, with a constant $c>0$ depending only on $d$.
@@
First assume that $\Gamma$ is derived from a quaternion algebra. Consider the set $\{\mathrm{id},\sigma,\sigma_{3},\ldots,\sigma_{d}\}$ of the embeddings of $K$. We recall that $|\sigma_{j}(\mathrm{tr}(\gamma))|<2$ for any non-identity element $\gamma\in\Gamma^{(2)}$ and for all $j=3,\ldots,d$. Let $s\neq t\in(2,T]$ be two traces of elements of $\Gamma$. By inequality (4) from Lemma 3.1 applied to $|\sigma(s)|$ and $|\sigma(t)|$, we have
@@
with $C_{1}=4^{d-1}>0$. These inequalities show that $t$ and $s$ are $cT^{\delta-1}$ separated with $c=\frac{1}{C_{1}}>0.$
In general, similar to [LS94, Lemma 2.1], we consider all the embeddings of the field $L=\mathbb{Q}(\mathcal{L}(\Gamma))$ into $\mathbb{C}$ which we denote by $\tau_{1}=\mathrm{id},\tau_{2},\ldots,\tau_{n}$, $n=[L:\mathbb{Q}]$. We order them so that the first $k$ embeddings extend the embedding $\sigma_{1}=\mathrm{id}$ of the subfield $K$ of $L$, and the next $k$ embeddings extend $\sigma_{2}=\sigma$. Since $\mathrm{tr}(\gamma)^{2}=\mathrm{tr}(\gamma^{2})+2$ we have $x^{2}-2\in\mathcal{L}(\Gamma^{(2)})$ for any $x\in\mathcal{L}(\Gamma)$, in particular, $x^{2}\in K$.
Suppose there are more than $2^{2k-1}(M+1)$ points of $\mathcal{L}(\Gamma)$ in the interval [N1,N][N-1,N], with $M$ a positive integer. Then by the pigeonhole principle there are at least two different elements $t,s\in\mathcal{L}(\Gamma)$ and some $\epsilon_{i}\in\{-1,1\}$, $i=2,\ldots,2k$, such that $0<|t-s|\leq\frac{1}{M}$, $\tau_{i}(t)=\epsilon_{i}t$, $\tau_{i}(s)=\epsilon_{i}s$ for $2\leq i\leq k,$ and $\tau_{i}(t)=\epsilon_{i}\sqrt{\sigma_{2}(t^{2})}$, $\tau_{i}(s)=\epsilon_{i}\sqrt{\sigma_{2}(s^{2})}$ for $k<i\leq 2k$.
Let $0\neq t-s\in\mathcal{O}_{L}$ be two such elements. For $1\leq i\leq k$, we have $|\tau_{i}(t-s)|\leq\frac{1}{M}$ by the assumption. For the next $k$ embeddings $k<i\leq 2k$, $|\tau_{i}(t-s)|\leq 4N^{1-\delta}$ by Lemma 3.1 applied to $\gamma^{2}$ and our assumption on $\tau_{i}$. And for the remaining for $i>2k$, we have $|\tau_{i}(t-s)|\leq 4$ (we assume that $\tau_{i}(\mathcal{L}(\Gamma))\subset[-2,2]$ for $i>2k$, which follows from $\mathrm{tr}(\gamma)^{2}=\mathrm{tr}(\gamma^{2})+2$ and $\sigma_{j}(\mathcal{L}(\Gamma^{(2)}))\subset[-2,2]$, for $j>2$). Hence we show that
$1\leq|\mathrm{N}_{L\mid\mathbb{Q}}(t-s)|\leq\frac{1}{M^{k}}(4N)^{(1-\delta)k}4^{n-2k},$
which implies that $M\leq CN^{1-\delta}$ with $C=C(\Gamma)>0$. ∎
@@
There exist $0<\delta<1$ and $C>0$ depending on $\Gamma$ such that for $T>2$,
$(\mathcal{L}(\Gamma)\cap[2,T])<CT^{2-\delta}.$
Proof.
We have
$(\mathcal{L}(\Gamma)\cap[2,T])\leq\sum_{N=3}^{T}\#(\mathcal{L}(\Gamma)\cap[N-1,N])<CT^{1-\delta}T.$
@@
First assume that $\Gamma$ is a non-arithmetic group as above.
Let $\mathcal{N}^{\prime}(\ell)$ be the number of distinct lengths bounded by $\ell$ of closed geodesics in the orbifold $\Gamma\backslash\mathbb{H}$. Each length produces a unique positive trace of a hyperbolic element of $\Gamma$ bounded by $2\cosh({\ell}/{2})$, hence by Corollary 3.3 we have
(7) $\mathcal{N}^{\prime}(\ell)\leq C\left(2\cosh\left(\frac{\ell}{2}\right)\right)^{2-\delta}.$
Since $\cosh(\frac{\ell}{2})\leq e^{\frac{\ell}{2}}$, we obtain
@@
for some constants $C^{\prime}>0$ and $0<\delta^{\prime}<\frac{1}{2}.$ By the prime geodesic theorem (cf. [Sar80, Theorem 3.4]), the number $\mathcal{N}(\ell)$ of closed geodesics of $\Gamma\backslash\mathbb{H}$ counted with multiplicity is given by
$\mathcal{N}(\ell)\sim\dfrac{e^{\ell}}{\ell}$ when $\text${ when $\ell\to\infty$.
By differentiating the integral formula for $\langle g(\ell)\rangle$ we see that the asymptotic behavior of the mean multiplicity is determined by the average derivatives of the counting functions (cf. [Bol93b, Section 3.2]):
$\dfrac{\mathcal{N}(\ell)}{\ell}\sim\langle g(\ell)\rangle\dfrac{\mathcal{N^{\prime}}(\ell)}{\ell}\text{, when }\ell\to\infty.$
Thus the mean multiplicity of closed geodesics of length at most $\ell$ grows as
$\langle g(\ell)\rangle\geq c\dfrac{e^{\delta^{\prime}\ell}}{\ell}$
for sufficiently large $\ell$, where the constants $c>0$ and $0<\delta^{\prime}<\frac{1}{2}$ depend on $\Gamma$.
If $\Gamma$ is arithmetic the result is well known. It follows from the bounded clustering property of Luo and Sarnak [LS94, Lemma 2.1] applied the same way that we use our Lemma 3.2 (see also [Bol93a] for more details). ∎
@@
Since $X$ has curvature $\leq\kappa$ for some $\kappa\leq 0$ and $S$ has curvature $\leq-1$ in the sense of Alexandrov, we can apply the generalized Kirszbraun theorem [LS97, Theorem A] in order to obtain an extension $g$ of $f\mid\Omega$ with the same Lipschitz constant. ∎
@@
Let $\Gamma$ be a non-cocompact semi-arithmetic Fuchsian group of arithmetic dimension $2$ which admits a modular embedding. Then there exists a constant $0<\delta<1$ depending on $\Gamma$ such that
$|\sigma(\mathrm{tr}(\gamma))|\leq 2|\mathrm{tr}(\gamma)|^{1-\delta}$
@@
[…]
$G(\Gamma z)=\Delta(z,F(z)).$